OpenStudy (anonymous):

simplify (2/5)^2... HOW???

OpenStudy (anonymous):

4/25

OpenStudy (across):

\[\left(\frac{2}{5}\right)^2=\frac{2^2}{5^2}=\frac{4}{25}\]

OpenStudy (karatechopper):

whenever you are put to a power its the number times itself however many times the exponent says

OpenStudy (anonymous):

Thanks